OpenStudy (anonymous):

Solve for u. 7u+19=2(u-3) Simplify your answer as much as possible.

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

7u+19 = 2u-6 7u-2u = -6-19 5u = -25 u= -5

OpenStudy (anonymous):

solve for v. -7(v+9)=-3v-27

OpenStudy (anonymous):

-7v-63 = 3v-27 -7v-63= 3v-27 -7v-3v= -27+63 -10v= 36 v=- 36/10

OpenStudy (anonymous):

v = - 18/5

OpenStudy (anonymous):

so -3.6

OpenStudy (anonymous):

yup